OpenStudy (whateveriwant):

Which of the following is a function equivalent to f(x) = 3(1.7)4x? f(x) = 25.06x f(x) = 3(8.35)x f(x) = 3(8.25x) f(x) = 487.97x

OpenStudy (mathmale):

Welcome to OpenStudy! Caution: Your " f(x) = 3(1.7)4x" is vague. I believe you mean " f(x) = 3(1.7)^(4x)." Please use the " ^ " symbol to denote exponentiation. Can you think of a way to simplify (reduce to lowest terms) " f(x) = 3(1.7)^(4x)"?

OpenStudy (mathmale):

\[f(x) = 3(1.7)^{4x}\]

OpenStudy (whateveriwant):

I all ready figured it out but thank you!

OpenStudy (mathmale):

Please re-write the following: f(x) = 25.06x f(x) = 3(8.35)x f(x) = 3(8.25x) f(x) = 487.97x using the " ^ " to denote exponentiation.

OpenStudy (mathmale):

You're welcome.